The height of a triangle is 1cm less than the length of the base. The area is 1770cm squared. Find the height.
Answer by checkley77(12569)

A=BH/2
H=B-1
1770=B(B-1)/2
1770*2=B^2-B
B^2-B=3,540
B^2-B-3,540=0
(B-60)(B+59)=0
B-60=0
B=60 CM.
H=60-1=59 CM. ANS. FOR THE HEIGHT.
PROOF:
1770=60*59/2
1,770=3,540/2
1,770=1,770